### v3

# GEMINUS: Dual-aware Global and Scene-Adaptive Mixture-of-Experts for End-to-End Autonomous Driving

Abstract
End-to-end autonomous driving requires adaptive and robust handling of complex and diverse traffic environments. However, prevalent single-mode planning methods attempt to learn an overall policy while struggling to acquire diversified driving skills to handle diverse scenarios. Therefore, this paper proposes GEMINUS, a Mixture-of-Experts end-to-end autonomous driving framework featuring a Global Expert, a Scene-Adaptive Experts Group, and equipped with a Dual-aware Router. Specifically, the Global Expert is trained on the overall dataset, possessing robust performance. The Scene-Adaptive Experts are trained on corresponding scene subsets, achieving adaptive performance. The Dual-aware Router simultaneously considers scenario-level features and routing uncertainty to dynamically activate expert modules. Through the effective coupling of the Global Expert and the Scene-Adaptive Experts Group via the Dual-aware Router, GEMINUS achieves adaptive and robust performance in diverse scenarios. GEMINUS outperforms existing methods in the Bench2Drive closed-loop benchmark and achieves state-of-the-art performance in Driving Score and Success Rate, even with only monocular vision input. Furthermore, ablation studies demonstrate significant improvements over the original single-expert baseline: 7.67% in Driving Score, 22.06% in Success Rate, and 19.41% in MultiAbility-Mean. The code will be available at https://github.com/newbrains1/GEMINUS.
Figure 1: The overall architecture of GEMINUS. GEMINUS integrates a Global Expert and a Scene-Adaptive Experts Group. During the training stage, the Global Expert is trained on the overall dataset. Concurrently, each scene-adaptive expert is trained on its respective scene subset guided by scenario-aware routing. In the inference stage, as shown above, features extracted from upstream encoders are processed by the Dual-aware Routing. When scenario uncertainty is below a threshold, the scenario-aware routing activates the highest-scoring adaptive expert; conversely, for high uncertainty, the uncertainty-aware routing activates the Global Expert. This design ensures both adaptive and robust performance across diverse scenarios.
I INTRODUCTION
In recent years, a prominent research direction in autonomous driving has been the development of planning-oriented end-to-end models[1]. In contrast to modular autonomous driving consisting of modular pipelines such as perception, prediction and planning [2, 3, 4], end-to-end methods directly map raw sensor inputs to planned trajectories [5, 6, 7, 8], control signals [9, 10], or a fused output derived from trajectory and control branches [11, 12]. These approaches provide a holistic model for driving, enabling unified optimization towards a global objective, significantly reducing manual engineering efforts, and allowing for the direct use of rich sensor information.
Despite their notable benefits, a persistent limitation of current end-to-end autonomous driving models stems from their global imitation learning on overall training datasets. This approach, typically employing single-mode planning with L2 loss, inherently models the complex output space as a single Gaussian distribution, leading to a tendency towards mode averaging[8, 13]. Consequently, their performance is compromised, as the generated output represents an averaged behavior across diverse scenarios rather than the optimal policy for the current specific scenario. This ultimately restricts the acquisition of diversified driving skills to handle diverse scenarios. Prior approaches employed command-based conditional imitation learning to mitigate mode averaging[9, 14]. However, this approach faced an inherent limitation: solely relying on driving commands is insufficient to distinguish complex scenarios (e.g., an overtaking scenario simultaneously includes turn left, go straight, and turn right commands). Such rigid classification fails to comprehensively consider rich scene information, thus hindering the capture of the diversity of driving skills.
Inspired by the success of Mixture-of-Experts (MoE) architectures in large language models (LLMs) to handle complex data distribution[15], MoE architectures present significant potential for addressing challenges in autonomous driving. By offering a fine-grained scenario adaptation and specialized behavior generation, MoE could mitigate the mode averaging problem and enhance model adaptability in diverse driving scenarios. However, directly transferring generic MoE architectures, designed primarily for static textual data, to autonomous driving reveals an inherent unsuitability. Specifically, they struggle with effective expert specialization due to lack of explicit scenario division, and fail to adequately consider the robustness requirements of autonomous driving.
Therefore, this paper proposes GEMINUS: dual-aware Global and scEne-adaptive MIxture of experts for end-to-end autoNomoUS driving. Specifically, the Global Expert is trained on the overall dataset, possessing robust performance. The Scene-Adaptive Experts are trained on corresponding scene subsets, achieving adaptive performance. The Dual-aware Router simultaneously considers scenario-level features and routing uncertainty to dynamically activate expert modules. Through the effective coupling of the Global Expert and the Scene-Adaptive Experts Group via the Dual-aware Router, GEMINUS simultaneously achieves adaptive and robust performance in diverse scenarios. Our contributions can be summarized as follows:
  • GEMINUS is proposed as a novel Mixture-of-Experts end-to-end autonomous driving framework. This framework effectively integrates a Global Expert and a Scene-Adaptive Experts Group via Dual-aware Router, to simultaneously achieve adaptive performance in feature-distinct scenarios and robust performance in feature-ambiguous scenarios.
  • This paper introduces a Dual-aware Router for end-to-end autonomous driving, uniquely designed with scenario-awareness to identify differences between diverse scenarios, and uncertainty-awareness to model routing uncertainty.
  • To further investigate GEMINUS’s intrinsic routing mechanism, we analyze the impact of the uncertainty threshold on driving performance and examined the router accuracy and expert utilization on the validation set.
II RELATED WORK
II-A End-to-End Autonomous Driving
End-to-end autonomous driving is an emerging paradigm. It directly maps raw sensor inputs to vehicle control actions or planned trajectories. This approach offers significant advantages by simplifying system architecture and mitigating error propagation inherent in cascaded modular pipelines. End-to-end driving policies are predominantly learned using Imitation Learning (IL). This typically involves behavior cloning (BC) to mimic expert demonstrations and capture human-like driving behaviors [5, 6, 7, 8, 9, 10, 11, 12]. Reinforcement Learning (RL) also plays a role. It enables dynamic policy optimization through environmental interaction and reward design [16, 17, 18].
In 2019, CILRS[14] was proposed to obtain control signals by introducing conditional imitation learning. Trajectory-guided Control Prediction (TCP) [11] is a simple and robust monocular vision baseline. This method innovatively integrates trajectory planning and direct control into a unified pipeline for joint learning and predictive fusion. Furthermore, techniques such as TransFuser [5] adopt the Transformer to fuse information from heterogeneous sensors (cameras and LiDAR). Additionally, DriveAdapter [10] employs a student model to learn rich environmental representations from multi-camera information, aiming to overcome the traditional coupling barriers between perception and planning. Beyond fusion, recent innovations also include vectorized scene representations, exemplified by VAD [19], which models driving scenes as fully vectorized elements to improve planning efficiency and robustness. Hydra-MDP [20] explores multimodal planning by distilling knowledge from human and rule-based teachers to generate diverse trajectory candidates. Diffusion models have also emerged as a powerful tool, with DiffusionDrive [21] utilizing truncated diffusion policies to model multi-modal action distributions while simultaneously achieving real-time control. Despite these advancements, existing end-to-end models remain constrained by mode averaging, making it challenging to effectively handle diverse scenarios.
II-B Mixture-of-Experts in Autonomous Driving
MoE architecture has emerged as a significant method for scaling large language models and enhancing task specialization. In Large Language Models, sparse MoE designs boost model capacity and processing efficiency through conditional computation [15]. The strength of MoE lies in its ability to leverage individual experts’ strengths across varying data subsets or tasks, thereby improving overall model performance. In [22], a task-level MoE was applied to multilingual translation, intelligently routing inputs based on linguistic or task identifiers to achieve performance gains and improved inference throughput.
Despite promising results in LLMs, MoE’s application in end-to-end autonomous driving remains underexplored. Some existing studies have explored MoE architectures in autonomous driving for tasks like rare scenario perception [23], long-tailed trajectory prediction[24], domain adaption in different weathers [25], safe trajectory prediction and planning [26], and facilitating generalization of planner [27]. However, these existing approaches have not focused on leveraging MoE to enhance adaptive and robust performance in diverse scenarios.
III METHODOLOGY
Fig. 1 illustrates the overall architecture of GEMINUS end-to-end autonomous driving framework. Informed by certain design philosophies of TCP [11], a single-expert baseline is established. Building upon this baseline, we integrate it with Dual-aware MoE, culminating in GEMINUS end-to-end autonomous driving framework.
III-A Preliminaries
End-to-End Autonomous Driving. The objective of end-to-end autonomous driving is to directly map raw sensor inputs to corresponding trajectories or control actions. In this paper, the raw sensor input $x$ encompasses: a front-facing camera image $i$, the ego-vehicle speed $v$, a high-level navigation command $c$, and a goal point $(x_{g},y_{g})$. The raw sensor inputs are processed by the end-to-end model. Encoders first process these inputs to generate intermediate features. These features are then fed into a trajectory planner. The trajectory planner generates a planned trajectory, comprising waypoints over $K$ steps. This planned trajectory is then fed into a Proportional-Integral-Derivative (PID) controller. The controller subsequently produces the final longitudinal control signals: throttle $\in[0,1]$, brake $\in[0,1]$, and the lateral control signal: steer $\in[0,1]$.
Mixture-of-Experts. MoE architectures offer a principled approach to address the complexities of multimodal data distributions by employing a “divide and conquer” strategy [15]. Introducing the MoE Framework to end-to-end trajectory planning, the overall policy distribution $p_{\theta}(Y\mid X)$ is typically represented as a probabilistic mixture of policy distributions of components $K$, each parameterized by an expert $m_{\theta}(Y\mid Z=k,X)$ and weighted by a gating network $q_{\theta}(Z=k\mid X)$, formalized as:
$p_{\theta}(Y\mid X)=\sum_{k=1}^{K}q_{\theta}(Z=k\mid X)\cdot m_{\theta}(Y\mid Z=k,X)$ (1)
By flexibly coupling multiple policy distributions, this probabilistic formulation offers a promising framework to effectively model multimodal distributions in end-to-end trajectory planning and tackle the mode averaging problem. Despite potential non-convex optimization challenges in learning such mixture models, deep learning implementations of MoE often simplify this by identifying and assigning the most suitable expert. Notably, models employing a Hard Assignment approach (i.e., selecting a single ”best” expert for a given sample) are highly effective and computationally efficient for multimodal distributions[28]. This is because Hard Assignment directly selects the most matching behavior mode, avoiding the averaging of all experts’ outputs, which further mitigates the mode averaging problem. Building upon these theoretical underpinnings, this paper proposes GEMINUS, a distinctive MoE framework that is specifically tailored for diverse and complex autonomous driving scenarios. At its core, a Dual-aware Router possesses scenario and uncertainty awareness to dynamically activate experts from a Global Expert and a Scene-Adaptive Experts Group. During inference, the Dual-aware Router processes intermediate features $x$ extracted by the encoders. It determines the final output $y$ based on the uncertainty measure $U(x)$ and the scores of scene experts $S_{E}(x)$, formalized as:
$y=\begin{cases}f_{\text{global}}(x),&\text{if }U(x)\geq\tau\\ f_{\operatorname*{argmax}\limits_{i\in\mathcal{S}}\;S_{E}(x)}(x),&\text{if }U(x)<\tau\end{cases}$ (2)
Here $x$ is the input feature. $\tau$ denotes a predefined uncertainty threshold. $S$ is the set of all the scene-adaptive experts. When $U(x)$ is low ($U(x)<\tau$), the expert with the highest routing score $S_{E_{i}}(x)$ is selected. This achieves precise and scenario-specific planning in feature-distinct scenarios. In contrast, when the uncertainty of the scenario $U(x)$ is high ($U(x)\geq\tau$), the model selects the Global Expert $f_{\text{global}}(x)$. This ensures robust performance in feature-ambiguous scenarios. Such design allows GEMINUS to effectively avoid the mode averaging problem, thereby achieving adaptive and robust performance in diverse scenarios.
III-B Single-Expert Baseline
Feature Encoders. The image encoder is built on a ResNet34 architecture, pre-trained on ImageNet [29]. This encoder processes the front-facing camera input image and outputs a 1000-dimensional feature embedding vector $I_{\text{feature}}$. Concurrently, a measurement encoder receives a concatenated input $m$ and generates a 128-dimensional measurement feature vector $M_{\text{feature}}$. The input $m$ comprises ego-vehicle speed $v$, a high-level navigation command $c$, and the navigation goal point $(x_{g},y_{g})$. Finally, $I_{\text{feature}}$ and $M_{\text{feature}}$ are concatenated to form combined feature $F$ for the subsequent trajectory planner and router.
Trajectory Planner. The trajectory planner receives the upstream combined feature vector $F$ as input. This input is then passed through a series of linear layers for downsampling, forming a 256-dimensional feature vector $f$. This feature vector $f$ is then fed into a waypoint generator GRU[30]. The GRU model auto-regressively generates future waypoints one by one. This forms a sequence of waypoints $(w_{0},w_{1},...,w_{k})$ for the next $K$= 4 steps. The longitudinal and lateral controllers process these waypoints to generate the final longitudinal control signals (throttle $\in[0,1]$, brake $\in[0,1]$) and the lateral control signal (steer $\in[0,1]$).
Figure 2: Dual-aware MoE vs Vanilla MoE. A vanilla MoE typically tries to balance the load across its experts. In contrast, Dual-aware MoE leverages scenario subset IDs during training. This explicitly directs inputs from specific scenario subsets to their corresponding experts through a scenario-aware routing. During inference, the learned router dynamically activates the most appropriate expert based solely on the input scene features, thus obviating the need for explicit scenario ID. Furthermore, an uncertainty-aware routing is introduced to ensure the model’s robustness and stability. Specifically, when scenario features are ambiguous, preventing the router from effectively distinguishing the true underlying expert, the input is strategically routed to a global expert, ensuring robust and stable model performance.
III-C Scenario-aware Routing Mechanism
Vanilla MoE aims to balance expert usage across GPUs to utilize maximum benefit from features of the inputs. However, this leads to inefficient knowledge sharing among experts when dealing with heterogeneous input distributions. For example, driving policies for a Merging scenario differ significantly from an Emergency Brake scenario. To address this inefficiency and foster specialized knowledge, a scenario-aware routing mechanism is introduced. This mechanism draws inspiration from dataset-aware routing in [31].
Inspired by the scenario classification in Bench2Drive [32], five autonomous driving scenario categories are classified: Merging, Overtaking, Emergency Brake, Give Way, and Traffic Sign. During the training phase, the scenario-aware router is explicitly trained to route input feature vectors $F$ based on their corresponding scenario category.
Let $\mathcal{S}=\{s_{m}\}_{m=1}^{|\mathcal{S}|}$ be a set of predefined scenario categories. An input feature vector $x$ belongs to a scenario $s_{m}$ ($x\in s_{m}$). We define a mapping function $h:S\rightarrow E$. This function assigns each scenario category $s_{m}$ to a specific scene-adaptive expert $e_{i}\in E$. Here, $E$ denotes the group of scene-adaptive experts. To enforce this routing strategy, a router loss is designed as $L_{scenario}$. This loss is formulated as a cross-entropy loss. It is computed between the router’s predicted expert selection probabilities $p_{i}(x)$ (representing the probability of selecting expert $e_{i}$) and the target expert label $h(s_{m})$ corresponding to input $x$ from scenario $s_{m}$:
$\mathcal{L}_{\text{scenario}}=-\sum_{i=1}^{|E|}\mathbb{1}(h(s_{m})=i)\cdot\log p_{i}(x)$ (3)
$1(\cdot)$ is the indicator function. This loss ensures that all inputs originating from a specific scenario category are primarily dispatched to their designated scene-adaptive expert. By selectively routing inputs based on their scenario feature, this mechanism promotes efficient knowledge specialization within each expert. This enables the model to learn adaptive driving policies.
III-D Uncertainty-aware Routing Mechanism
While scene-adaptive experts excel in feature-distinct scenarios, relying solely on them can be problematic in feature-ambiguous scenarios. This compromises robustness, especially in safety-critical applications like autonomous driving. To mitigate this risk and ensure reliable performance in diverse scenarios, an uncertainty-aware routing mechanism is introduced.
The raw input $x$ is first processed by encoders to form the feature vector $F$. The router then computes a probability distribution over experts from the feature vector $F$, denoted as $P(x)=[p_{1},p_{2},...,p_{N}]$. Subsequently, the Information Entropy[33] of this distribution is calculated to reflect the uncertainty of the router’s decision:
$H(P(x))=-\sum_{i=1}^{N}p_{i}\log(p_{i})$ (4)
To normalize this entropy to a $[0,1]$ range, it is divided by the theoretical maximum entropy. This maximum occurs when probabilities are uniformly distributed across all experts ( $i.e.,pi=1/N\,for\,all\,i$), and its value is $log(N)$. Thus, the Normalized Information Entropy $U(x)$ is defined as:
$U(x)=\frac{H(P(x))}{\log(N)}=\frac{-\sum_{i=1}^{N}p_{i}\log(p_{i})}{\log(N)}$ (5)
This Normalized Information Entropy $U(x)$ serves as the measure of scenario uncertainty. A value close to 0 indicates high certainty, meaning the scenario is distinct and the router is confident. Conversely, a value close to 1 indicates high uncertainty meaning the scenario is ambiguous and the router is undecided.
III-E Loss Design
GEMINUS is trained using a comprehensive loss function that combines multiple objectives.
Global Expert Loss. The Global Expert aims to provide a robust, generalized driving policy. Its loss $\mathcal{L}_{Global}$ comprises three main terms.
Trajectory Imitation Loss. This term encourages the Global Expert to accurately predict future waypoints. It minimizes the L1 distance between predicted and ground truth waypoints, formalized as:
$\mathcal{L}_{\text{traj\_global}}=\sum_{t=1}^{K}\left\|w_{t}-\hat{w}_{t}\right\|_{1}$ (6)
Where $w_{t}$ and $\hat{w}_{t}$ are the ground truth and predicted waypoints, respectively, at step $t$ within a prediction horizon of $K$ steps.
Feature Alignment Loss. This loss ensures consistent feature representation. It measures the L2 distance between the Global Expert’s output features and the corresponding expert feature. This serves as an additional supervision signal [18], formalized as:
$\mathcal{L}_{F\_global}=\left\|j_{\text{global}}-j_{\text{expert}}\right\|_{2}$ (7)
The $j_{global}$ denotes the intermediate feature representation from the Global Expert, and $j_{expert}$ is the corresponding feature from the expert for alignment.
Value Alignment Loss. This term guides the Global Expert to predict the expected return for the current state. It employs an L2 loss, formalized as:
$\mathcal{L}_{V\_global}=\left\|v_{\text{global}}-v_{\text{expert}}\right\|_{2}^{2}$ (8)
The $v_{global}$ is the value predicted by the Global Expert’s value branch, and $v_{expert}$ is the corresponding value from Think2Drive[18] expert.
Global Expert Loss is combined as:
$\mathcal{L}_{\text{Global}}=\lambda_{\text{traj}}\,\mathcal{L}_{\text{traj\_global}}+\lambda_{F}\,\mathcal{L}_{F\_global}+\lambda_{V}\,\mathcal{L}_{V\_global}$ (9)
$\lambda_{traj}$, $\lambda_{F}$, and $\lambda_{V}$ are tunable loss weights.
Scene-Adaptive Experts Group Loss. The Scene-Adaptive Experts Group comprises $N$ distinct experts. Each expert is trained to master policies for specific scenarios. The loss for this group, $L_{Adaptive}$, is computed as a weighted sum of individual expert losses. For a given sample $x$, only the adaptive expert it is routed to contributes to the loss. If $x$ is routed to expert $e_{i}$, its loss is calculated. This calculation is similar to the components of the Global Expert Loss. It is formalized as:
$\begin{split}\mathcal{L}_{\text{Adaptive}}&=\mathbb{1}(x\rightarrow e_{k})\cdot\Big{(}\lambda_{\text{traj}}\,\mathcal{L}_{\text{traj},k}(x)\\ &\quad+\lambda_{F}\,\mathcal{L}_{F,k}(x)+\lambda_{V}\,\mathcal{L}_{V,k}(x)\Big{)}\end{split}$ (10)
For integer $k\in{1,...,N}$, $\mathcal{L}_{traj,k}(x)$, $\mathcal{L}_{F,k}(x)$, and $\mathcal{L}_{V,k}(x)$ are the trajectory imitation, feature alignment, and value prediction losses for expert $e_{k}$ on sample $x$. These are similar to those defined for the Global Expert. $1(\cdot)$ is the indicator function, ensuring that only the activated expert’s loss contributes for that specific sample.
Router Loss. The Router Loss is designed to effectively train the Dual-aware Router to make accurate expert selection decisions. This loss corresponds to $\mathcal{L}_{\text{scenario}}$, as described in Equation (3).
Speed Prediction Loss. To enhance the agent’s ability to estimate its current dynamic state, a dedicated speed prediction head is integrated. This head predicts the current ego-vehicle speed from the encoded feature. An L1 loss is employed for this prediction. It minimizes the absolute difference between the predicted and ground truth speeds, denoted as $L_{speed}$.
Total Loss. The total loss function for training the GEMINUS model is the weighted sum of all aforementioned loss components:
$\begin{split}\mathcal{L}_{\text{total}}&=\lambda_{\text{Global}}\,\mathcal{L}_{\text{Global}}+\lambda_{\text{Adaptive}}\,\mathcal{L}_{\text{Adaptive}}\\ &\quad+\lambda_{\text{scenario}}\,\mathcal{L}_{\text{scenario}}+\lambda_{\text{speed}}\,\mathcal{L}_{\text{speed}}\end{split}$ (11)
The $\lambda_{Global}$, $\lambda_{Adaptive}$, $\lambda_{scenario}$, and $\lambda_{speed}$ are empirically determined weighting coefficients. They balance the contributions of each loss term.
TABLE I: Closed-Loop and Open-Loop Performance on Bench2Drive Benchmark.
Method Venue Input Open-loop Metric Closed-loop Metric
Avg. L2 (m)$\downarrow$ Driving Score $\uparrow$ Success Rate(%) $\uparrow$
TCP* [11] NeurIPS 2022 Ego State + Front Camera $1.70$ $40.70$ $15.00$
TCP-ctrl* [11] NeurIPS 2022 Ego State + Front Camera $30.47$ $7.27$
TCP-traj* [11] NeurIPS 2022 Ego State + Front Camera $1.70$ $59.90$ $30.00$
UniAd-Base [6] CVPR 2023 Ego State + 6 Cameras $0.73$ $45.81$ $16.36$
ThinkTwice* [35] CVPR 2023 Ego State + 6 Cameras $0.95$ $62.44$ $31.23$
VAD [7] ICCV 2023 Ego State + 6 Cameras $0.91$ $42.35$ $15.00$
DriveAdapter* [10] ICCV 2023 Ego State + 6 Cameras $1.01$ $64.22$ $33.08$
GenAD [36] ECCV 2024 Ego State + 6 Cameras $44.81$ $15.90$
DriveTrans [8] ICLR 2025 Ego State + 6 Cameras 0.62 $63.46$ $35.01$
MomAD [37] CVPR 2025 Ego State + 6 Cameras $0.82$ $47.91$ $18.11$
SparseDrive [38] ICRA 2025 Ego State + 6 Cameras $0.83$ $42.12$ $15.00$
TTOG [39] ArXiv 2025 Ego State + 6 Cameras $0.74$ $45.23$ $16.36$
GEMINUS* Ego State + Front Camera $1.60$ 65.39 37.73
Avg. L2 is averaged over the predictions in 2 seconds under 2Hz. * denotes expert feature distillation.
TABLE II: MultiAbility Results on Bench2Drive Benchmark.
Method Venue Input Ability (%$\uparrow$
Merging Overtaking Em-Brake Give Way Traffic Sign Mean
TCP* [11] NeurIPS 2022 Ego State + Front Camera $16.18$ $20.00$ $20.00$ $10.00$ $6.99$ $14.63$
TCP-ctrl* [11] NeurIPS 2022 Ego State + Front Camera $10.29$ $4.44$ $10.00$ $10.00$ $6.45$ $8.23$
TCP-traj* [11] NeurIPS 2022 Ego State + Front Camera $8.89$ $24.29$ $51.67$ $40.00$ $46.28$ $34.22$
UniAd-Base [6] CVPR 2023 Ego State + 6 Cameras $14.10$ $17.78$ $21.67$ $10.00$ $14.21$ $15.55$
ThinkTwice* [35] CVPR 2023 Ego State + 6 Cameras $27.38$ $18.42$ $35.82$ 50.00 $54.23$ $37.17$
VAD [7] ICCV 2023 Ego State + 6 Cameras $8.11$ $24.44$ $18.64$ $20.00$ $19.15$ $18.07$
DriveAdapter* [10] ICCV 2023 Ego State + 6 Cameras 28.82 $26.38$ $48.76$ 50.00 56.43 42.08
DriveTrans [8] ICLR 2025 Ego State + 6 Cameras $17.57$ $35.00$ $48.36$ $40.00$ $52.10$ $38.60$
SparseDrive [38] ICRA 2025 Ego State + 6 Cameras $12.50$ $17.50$ $20.00$ $20.00$ $23.03$ $18.60$
TTOG [39] ArXiv 2025 Ego State + 6 Cameras $16.18$ $24.29$ $20.00$ $21.50$ $23.03$ $21.12$
GEMINUS* Ego State + Front Camera $11.11$ 37.50 55.00 $40.00$ $45.26$ $37.77$
* denotes expert feature distillation.
TABLE III: Ablation Study on Bench2Drive Benchmark
Method DrivingScore SuccessRate MultiAbility
GEMINUS 65.39 37.73 37.77
ScenarioMoE-E2E (w/o 1) $62.38$ $32.27$ $34.46$
VanillaMoE-E2E (w/o 1+ 2) $59.23$ $29.09$ $32.05$
SingleExpert-E2E (w/o 1+ 2+ 3) $60.73$ $30.91$ $31.63$
For consistency, VanillaMoE-E2E employs five experts with Top-1 sparse activation. 1 denotes uncertainty-aware routing and Global Expert. 2 denotes scenario-aware routing. 3 denotes Mixture-of-Experts.
TABLE IV: Router Accuracy In Different Scenarios
Scenario Overall Merging Overtaking Em-Brake Give Way Traffic Sign
Accuracy 68.06% 32.85% 91.35% 54.03% 2.87% 90.45%
TABLE V: Expert Utilization In Different Scenarios
Expert Expert Utilization (%)
Overall Merging Overtaking Emergency Brake Give Way Traffic Sign
Global Expert $6.29$ $6.43$ $1.09$ $10.52$ $6.70$ $6.04$
Merging Expert $8.44$ 32.37 $0.16$ $3.51$ $4.07$ $2.62$
Overtaking Expert $19.22$ $3.13$ 91.04 $1.75$ $61.72$ $0.84$
Em-Brake Expert $16.08$ $3.67$ $5.66$ 52.33 $15.07$ $5.03$
Give Way Expert $0.23$ $0.00$ $0.31$ $0.24$ 2.15 $0.17$
Traffic Sign Expert $49.73$ $54.40$ $1.74$ $31.65$ $10.29$ 85.30
IV EXPERIMENTS
IV-A Experimental Setup
Dataset. GEMINUS is trained on the official Bench2Drive training dataset. This dataset is collected by Think2Drive [18], a reinforcement learning expert with latent world model. To ensure a fair comparison with existing baselines, this paper utilizes the base dataset (1000 clips) for training and open-loop validation. This dataset comprises a 950-clip training set and a 50-clip open-loop validation set. Each clip represents a specific traffic scenario, spanning approximately 150 meters.
Evaluation Metrics. For closed-loop evaluation, GEMINUS is assessed on 220 routes officially provided by Bench2Drive [32]. These short routes are structured into 44 interactive scenarios, with 5 distinct routes per scenario. The closed-loop evaluation metrics include a Driving Score, Success Rate, and five MultiAbility metrics defined by Bench2Drive: Merging, Overtaking, Emergency Brake, Give Way, and Traffic Sign.
Implementation Details. The resolution of the input RGB image is 900 $\times$ 256 pixels. The future prediction steps are set to $K$ = 4, with a prediction frequency of 2 Hz. For PID control settings, we adopt the well-tuned parameters proposed in Transfuser [4]. Specifically, for longitudinal control, the PID parameters are set to $K_{P}$ = 5.0, $K_{I}$ = 0.5, and $K_{D}$ =1.0. For lateral control, the PID parameters are $K_{P}$ = 0.75, $K_{I}$ = 0.75, and $K_{D}$ = 0.3. As for the uncertainty threshold $\tau$, the optimal value $\tau$ = 0.5 is determined through experiments. During the training phase, the training dataset is initially divided into five major scenario subsets. This division is inspired by Bench2Drive’s classification of driving skills. Each subset contains samples with specific scenario ID. For the Merging, Overtaking, Emergency Brake, Give Way, and Traffic Sign subsets, the respective scenario IDs are $[0,1,2,3,4]$. The loss function weight coefficients are configured as follows: $\lambda_{traj}$ = 1, $\lambda_{F}$ = 0.05, $\lambda_{V}$ = 0.001, $\lambda_{Global}$ = 1, $\lambda_{Adaptive}$ = 1, $\lambda_{scenario}$ = 1, and $\lambda_{speed}$ = 0.05. For all experiments, the model is trained on a single NVIDIA GeForce RTX 4090 GPU using a batch size of 96 for 32 epochs. The Adam optimizer is employed with an initial learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-7}$. The learning rate is reduced by a factor of 2 after 30 epochs.
IV-B Comparison with State-of-the-Art works
As shown in Table I, GEMINUS achieves state-of-the-art performance on Bench2Drive closed-loop benchmark for both Driving Score and Success Rate. Notably, GEMINUS relies solely on monocular visual input, and surpasses existing methods on the Bench2Drive benchmark that use 6-camera images inputs.
While GEMINUS does not exhibit superior performance in open-loop average L2 error, such metrics primarily indicate model convergence rather than reliably assessing real-world driving. In contrast, closed-loop metrics offer a more robust evaluation of actual driving capabilities, a point emphasized by previous research such as TransFuser++ [34] and Bench2Drive [32].
When focusing solely on monocular vision methods, GEMINUS significantly improves upon existing state-of-the-art monocular vision method–TCP-traj*[11]. GEMINUS achieves a 9.17% increase in Driving Score, a 25.77% increase in Success Rate, and a reduction of 5.88% in open-loop average L2 error. Furthermore, as shown in Table II, a 10.37% increase in MultiAbility-Mean.
IV-C Ablation Study
As shown in Table III, ablation study yields critical insights into the contribution of each GEMINUS component.
Comparing VanillaMoE-E2E with SingleExpert-E2E. It is obvious that directly introducing a generic MoE framework which is commonly used in LLMs into autonomous driving does not improve model performance. Without specific adaptation, it even leads to a slight decrease in Driving Score and Success Rate. This substantiates our hypothesis: end-to-end autonomous driving systems demand a more tailored MoE framework. Such a framework should specifically address the diverse and complex nature of real-world driving scenarios.
Comparing ScenarioMoE-E2E with SingleExpert-E2E. The scenario-aware routing mechanism comprehensively improves model performance. The Driving Score improved by 2.72%, Success Rate by 4.40%, and MultiAbility-Mean by 8.95%. The introduction of this mechanism not only enhances the model’s adaptive performance in diverse scenarios but also makes its routing logic more interpretable.
Comparing GEMINUS with ScenarioMoE-E2E. Further incorporating the uncertainty-aware routing mechanism and the Global Expert yields additional performance gains. The Driving Score improved by 4.83%, Success Rate by 22.06%, and MultiAbility-Mean by 19.41%. The integration of the uncertainty-aware routing mechanism and the Global Expert significantly enhances the model’s robustness and stability. This is particularly true in ambiguous scenarios where the router cannot confidently determine the current situation.
Figure 3: Driving Score variation trend with uncertainty threshold.
IV-D Analysis of Uncertainty Threshold
To investigate the impact of the uncertainty threshold $\tau$ on model performance, the uncertainty threshold $\tau$ is varied from 0.0 to 1.0 with a step size of 0.1, and conduct a series of closed-loop evaluations on the Bench2Drive Benchmark. As depicted in Fig.3, the model’s Driving Score and Success Rate show a trend of initial increase followed by a decrease as $\tau$ gradually increases, reaching its optimum at $\tau$ = 0.5. This indicates that when the router’s uncertainty is less than 0.5, the selection made by the scenario-aware routing is reliable, and the performance of the adaptive experts contributes to improved model performance. Conversely, when the router’s uncertainty is greater than or equal to 0.5, the scenario-aware routing cannot make a reliable decision, necessitating the intervention of the Global Expert to ensure robust and stable performance.
IV-E Router Accuracy and Expert Utilization
To better understand the intrinsic routing dynamics of the GEMINUS framework, analysis is conducted on the Bench2Drive open-loop validation set. This analysis focused on two key aspects during open-loop evaluation: router prediction accuracy and expert utilization. Router prediction accuracy is defined as the proportion of samples where the router correctly identifies the corresponding scenario. Expert utilization refers to the activation rates of both the Global Expert and the five Scene-Adaptive Experts.
Router Accuracy. As depicted in Table IV, the router’s overall scenario prediction accuracy reached 68.06%. It is worth noting that the Traffic Sign subset overlaps with both the Merging and Emergency Brake subsets. In such cases, a single sample might pertain to multiple scenarios. Therefore, the actual prediction accuracy could be even higher. This indicates that the scenario-aware routing can accurately determine the current scenario in most cases. However, it struggles with an accurate prediction in a minority of scenarios. A closer examination of the five validation set scenarios reveals that, in the Overtaking and Traffic Sign scenarios, the router exhibits the highest prediction accuracy. This is mainly because these scenarios have salient visual cues, such as obstacles or traffic signs. These cues significantly enhance the router’s ability to accurately predict the scenario. In contrast, the Give Way scenario presents the lowest prediction accuracy of 2.89%. This discrepancy stems from two primary factors. First, the Give Way subset constitutes only 3.16% of the training set and 4.00% of the validation set. This represents an inherent data imbalance within the official Bench2Drive dataset. Second, GEMINUS relies on monocular visual input. This constrains its ability to detect rear-approaching vehicles in Give Way scenarios, thereby impeding accurate scenario prediction.
Expert Utilization. As depicted in Table V, the “Overall” column reveals a Global Expert utilization rate of 6.29%. This indicates that GEMINUS primarily prioritizes routing to scene-adaptive experts in most instances. This allows it to leverage their scenario-specific capabilities. The Global Expert is mainly invoked only in highly ambiguous scenarios to ensure robust and stable performance. Furthermore, a comparative analysis of the “Global Expert” row in Table V with the router accuracy in Table IV shows a clear pattern. Global Expert utilization is minimal in scenarios with higher routing prediction accuracy, such as Overtaking (1.09%) and Traffic Sign (6.04%). Conversely, in the three scenarios characterized by lower routing prediction accuracy, the model exhibits increased Global Expert utilization. This helps to maintain robustness and stable performance.
V CONCLUSIONS
This paper presents GEMINUS, a novel Dual-aware Mixture-of-Experts framework tailored for end-to-end autonomous driving. Through the effective coupling of the Global Expert and Scene-Adaptive Experts Group via Dual-aware intelligent routing, GEMINUS simultaneously achieves adaptive performance in feature-distinct scenarios and robust performance in feature-ambiguous scenarios. Closed-loop evaluation is conducted on Bench2Drive, GEMINUS outperforms existing methods and achieves state-of-the-art performance in Driving Score and Success Rate, relying solely on monocular visual input. Furthermore, ablation studies demonstrate significant improvements over the original single-expert baseline: 7.67% in Driving Score, 22.06% in Success Rate, and 19.41% in MultiAbility-Mean. In addition, the impact of the uncertainty threshold on model performance is analyzed to determine its optimal value. Furthermore, an in-depth analysis of router accuracy and expert utilization provides insights into GEMINUS’s internal routing mechanism.
This study is limited by the use of monocular camera inputs. To enable the router to consider scene information more comprehensively, the exploration of Dual-aware routing with multi-camera input remains a promising direction for future research. Furthermore, a promising research direction is to replace GEMINUS’s expert networks with Low-rank Adaptation (LoRA) modules, providing a lightweight Mixture-of-Experts plugin that is highly effective for the efficient fine-tuning of pretrained models.